# Robustifying Long-term Human-Robot Collaboration through a Hierarchical and Multimodal Framework

### v1

@@ Abstract
Long-term Human-Robot Collaboration (HRC) is crucial for developing flexible manufacturing systems and for integrating companion robots into daily human environments over extended periods. However, sustaining such collaborations requires overcoming challenges such as accurately understanding human intentions, maintaining robustness in noisy and dynamic environments, and adapting to diverse user behaviors. This paper presents a novel multimodal and hierarchical framework to address these challenges, facilitating efficient and robust long-term HRC. In particular, the proposed multimodal framework integrates visual observations with speech commands, which enables intuitive, natural, and flexible interactions between humans and robots. Additionally, our hierarchical approach for human detection and intention prediction significantly enhances the system’s robustness, allowing robots to better understand human behaviors. The proactive understanding enables robots to take timely and appropriate actions based on predicted human intentions. We deploy the proposed multimodal hierarchical framework to the KINOVA GEN3 robot and conduct extensive user studies on real-world long-term HRC experiments. The results demonstrate that our approach effectively improves the system efficiency, flexibility, and adaptability in long-term HRC, showcasing the framework’s potential to significantly improve the way humans and robots work together.
Index Terms:
Human-robot collaboration; Multimodality; Intention prediction; Hierarchical planning; Assembly.
@@ I INTRODUCTION
Human-robot collaboration (HRC) has emerged as a pivotal area of research, particularly as robots become more prevalent in human-centric environments. Although frameworks and technologies supporting short-term HRC have advanced significantly[1, 2], long-term HRC is still not as well developed [3]. In industrial settings, utilizing long-term HRC can significantly enhance productivity and adaptability in complicated manufacturing processes [4]. In-home environments, robots acting as companions or elderly care agents can improve quality of life through long-term HRC [5]. However, achieving effective and robust long-term HRC presents several challenges: 1) Accurate Understanding of Human Intentions: Robots must accurately interpret human actions and intentions in complex and long-term tasks. 2) Robust to Noise: Systems must remain robust against environmental noise and unexpected disturbances. 3) Efficient Collaboration: Interactions should be optimized in completing time for varying contexts. 4) Adaptable to Different Users: Systems should adapt to individual users with diverse behaviors and preferences. .
Figure 1: Overview of the hierarchical task graph. The task graph represents the long-term HRC task, such as toy car assembly (as shown on the left). At the higher level (middle section), each node denotes a subtask. Each subtask consists of a sequence of collaborative actions between the human and the robot (as shown on the right).
@@
Addressing these challenges is essential for the success of long-term HRC in real-world applications. While existing research has made contributions in individual aspects of challenges, there remains a critical need for integrated solutions that address several challenges simultaneously. For instance, studies such as [1, 2, 6] explore adaptable HRC through human motion prediction. However, these works often overlook the robustness to disturbances and the seamless integration required for effective human-robot collaboration. Similarly, to enhance the understanding of human intentions, research by [7, 8] explores multimodal HRC frameworks; however, these frameworks often depend on additional wearable devices. This reliance on wearable devices can limit the generalizability of these systems across different interaction contexts and user populations. Furthermore, some previous works often fail to integrate these techniques into a unified framework and lack real-world demonstrations of complex long-term HRC tasks [9].
In this work, we focus on enhancing long-term HRC, characterized by long sequences of interactions. We utilize a hierarchical task graph to illustrate these long-term HRC tasks. The task graph used in our experiment is depicted in fig. 1. The aim for humans and robots is to progress from the initial node (at the bottom ) to the completion node (at the top), which signifies the end of the task, through a series of interactions. Accurate prediction of human intentions is critical for the robot to determine its collaborative controls effectively. From the task graph perspective, predicting human intentions equates to forecasting the subsequent node in the task graph. Given the long sequences of these interactions, long-term HRC tasks are prone to compounding errors; therefore, predictions must be precise, resilient to disruptions, and tailored to the diverse behaviors of individual humans. This capability ensures that the robot can proficiently plan and execute collaborative controls.
To address these challenges, we have developed a multimodal framework that integrates both visual and auditory signals to enhance the accuracy of intention prediction and current-position estimation within the task graph. The reliance on a single modality often fails to provide accurate predictions over complex, long-term tasks. Further improvements include a hierarchical design in both the human pose detection and intention prediction modules. This design strategy enables the system to initially predict the human’s actions and subsequently refine these predictions using the task graph to correct the detected actions and the corresponding node positions. This hierarchical approach not only increases the precision of intention understanding but also minimizes the impact of environmental noise and disturbances. By combining multimodal observations with hierarchical prediction, our proposed framework aims to achieve robust, efficient long-term HRC.
@@
    We present an effective framework for solving long-term human-robot collaboration tasks.
@@
    We introduce a hierarchical pose detection algorithm that effectively mitigates disturbances in multi-human environments, ensuring reliable human detection.
@@
    We validate our framework in real-world complex toy car assembly tasks. Through physical experiments and user studies, we demonstrate that our approach significantly improves the robustness, time efficiency, and flexibility of human-robot collaboration.
@@ II RELATED WORKS
Long-term Human-Robot Collaboration. Previous research on human-robot collaboration (HRC) has primarily focused on short-term tasks with long-horizon action sequences[10, 11], rather than fully addressing long-term, complex tasks. While these studies explore important aspects of task planning over extended action sequences [12, 13], there is limited work specifically targeting long-term human-robot collaboration that achieves comprehensive, goal-oriented tasks. Pirk et al.[10] successfully recognized long-term human sub-assembly actions but did not demonstrate human-robot collaboration in real-world scenarios. In this work, we implement our framework on a real-world long-term complex assembly task.
@@
Human Pose Detection and Intention Prediction. High-fidelity and robust human behavior prediction is a key component of achieving safe HRC [18, 19]. Several approaches utilize human detection models, such as OpenPose [20], to identify human key points, which are then used to predict human behaviors including intentions and/or trajectories [2, 6]. Reducing disturbances in human pose detection has been extensively studied using signal filtering techniques [21, 22]. Traditional filtering methods primarily mitigate temporal disturbances [23]. In contrast, we focus on spatial aspects and propose a novel hierarchical framework to reduce disturbances in HRC. Additionally, we implemented an online model adaptation algorithm to adapt the prediction models to accommodate different users’ behaviors [24].
@@ III PROBLEM FORMULATION
The Long-Term HRC Task. We consider a scenario where a human $A_{H}$ collaborates with a robot $A_{R}$, to achieve a set of predefined goals within a long-term task $\mathcal{G}$. The ultimate goal, $g$, is the completion of the task, represented by reaching the top of the hierarchical task graph $\mathcal{G}$, as shown in fig. 1. We define $\mathcal{G}$ as a Hierarchical and Multi-Agent Temporal Plan Graph (TPG) [25], where task sequences are structured within a directional graph. The nodes $\mathcal{N}$ represent individual sub-tasks, and the directed edges denote the precedence of these nodes.
Figure 2: Example of a temporal plan graph for each subtask. Note that, for a given subtask, there may be multiple plans that can achieve the same goal. In this figure, two different human-robot plans (action sequences) are illustrated to achieve the same subtask from the same initial state.
@@
Sub-task Configuration. At the highest level, $\mathcal{G}$ is the overarching long-term HRC task. Within this framework, each node corresponds to shorter-term HRC tasks, referred to here as sub-tasks. Each subtask $N_{i}\in\mathcal{N}$ is also organized as a multi-agent TPG, which represents both the precedence and collaborative dynamics between human and robot actions [26]. Each vertex at the lowest hierarchy level represents an action performed by a human or robot. Figure 1 illustrates a three-level structure of the task graph—from the multi-agent long-term task level to the multi-agent short-term sub-task level, down to the single-agent shortest-term action level. A further depiction of the lowest level temporal plan graph is shown in fig. 2, emphasizing that in HRC, the robot must predict human intentions, which involves foreseeing the subsequent node a human might conduct based on their behavior’s inherent property.
@@
The Objective Function. Let $E$ denote the environment in the HRC task, incorporating both humans and objects. Let $P$ denote the task progress, which tracks the completion of subtasks (or nodes within the task graph), and updates as the task advances. At timestep $t$, the human’s intention $x_{H}^{t}$ is influenced by the current state of the environment $E^{t}$ and task progress $P^{t}$. The robot predicts human intentions, $\hat{x}_{H}^{t}$, using its learnable model $\hat{f}_{H}$:
@@
The predicted intentions guide the robot in generating a motion sequence $q=[q^{1},\cdots,q^{n}]$. The objective for the robot is to find an efficient and robust motion sequence $q$ under safety and dynamics constraints. Let $J_{g}$ represent the task-specific cost of achieving the goal $g$, such as a tracking cost. To reflect time efficiency, we include the task completion time $T_{g}$, weighted by $\lambda$, in the objective function. Thus, the robot’s objective in HRC is formulated as:
@@
where $M_{R}(\cdot)$ represents the robot motion planning function. Accurate and robust prediction of human intentions, $\hat{x}_{H}^{t}$, is therefore pivotal to the process.
@@ IV PROPOSED FRAMEWORK
To address these challenges of long-term HRC tasks as discussed in section I, we propose a hierarchical and multimodal HRC framework, as illustrated in fig. 3. The framework includes two perception modules from different modalities (a visual perception module and an audio perception module), four modules for task planning (pose detection, intention prediction, speech recognition, and robot motion planner), and a robot motion controller.
@@
  • Planning. Pose Detection Module is responsible for identifying human poses and extracting key point trajectories. Its hierarchical design allows the framework to mitigate disturbances in complex, multi-human environments, ensuring more robust performance. Intention Prediction Module predicts human intentions based on the trajectories of human keypoints and task progress. Similarly, the hierarchical design of this module enables robots to respond to human actions more robustly. Speech Recognition Module recognizes speech commands for audio-based robot motion generation and corrects errors in robot motion caused by inaccuracies in human intention prediction. Robot Motion Planner Module outputs the high-level robot motions based on the fusion of predicted human intention, recognized human speech commands, and task progress. Figure 4 presents a detailed overview of the framework’s hierarchical planning modules.
  • Control. Controller Module receives the high-level motions from the planner module and translates them into low-level robot actions for execution.
Figure 4: The overview of the planning modules.
@@ IV-A Multimodal Perception
In human-to-human interactions, visual observations alone often fail to fully capture human intentions, necessitating the integration of verbal communication for comprehensive context. Applying this principle to long-term HRC tasks, we propose a multimodal framework, as depicted in fig. 3. This framework leverages visual observation to detect human poses and predict human intentions. However, visual data alone can be ambiguous, as the same human pose may indicate different intentions depending on the context. To resolve these ambiguities, we incorporate audio observations to add more contextual information. Speech commands enable humans to make clarifications and adjustments to the robot’s motions, informed by its visual interpretations. This integration yields a more comprehensive understanding of human intentions and enhances the robot’s adaptability to dynamic user demands. The multimodal approach offers two significant advantages: 1) By integrating visual and audio signals, the robot achieves a comprehensive perception of human intentions, allowing for more complex and long-term collaborations. 2) The inclusion of speech enables more natural and flexible interactions, making the system more user-friendly and accessible.
We employ mutual information analysis to demonstrate the effectiveness of our multimodal framework. Let $O_{V}$ denote the robot’s visual observation from the environment $E$, captured by the RGBD camera; and $O_{A}$ denote the robot’s audio observation, captured by the microphone. The robot’s performance in achieving the goal $g$ is enhanced by these informative observations. We consider the mutual information between the goal and observations:
$I(g;O_{V},O_{A})=H(g)-H(g|O_{V},O_{A})$ (4)
where $I(\cdot,\cdot)$ denotes mutual information and $H(\cdot)$ denotes entropy. By the nature of information fusion, we have:
$H(g|O_{V},O_{A})\leq\min\{H(g|O_{V}),H(g|O_{A})\}$ (5)
This inequality holds because the combined information from fused modalities $(O_{V},O_{A})$ is no less than the information from single modality $O_{V}$ or $O_{A}$ individually. We also provide the proof in section -A. Combining the above expressions, we find:
$I(g;O_{V},O_{A})$ $\geq\max\{H(g)-H(g|O_{V}),H(g)-H(g|O_{A})\}$
$=\max\{I(g;O_{V}),I(g;O_{A})\}$ (6)
Therefore, we conclude that the multimodal framework yields greater mutual information between the goal and the observations than a single-modal framework, significantly enhancing the robot’s capability to handle complex and long-term collaboration tasks.
IV-B Hierarchical Pose Detection
Effective long-term HRC critically depends on accurately predicting human intentions, which are deduced from the estimated human poses obtained from visual observations. However, visual observations can often be noisy, leading to errors in pose detection and subsequent intention prediction, ultimately impacting the robot’s motion control. This noise can arise from various sources, such as the presence of multiple humans in the scene or cluttered backgrounds. To mitigate these disturbances, we propose a hierarchical pose detection module that enhances both accuracy and efficiency. Our module operates through a two-step approach, based on the distance between the human and the robot, as outlined in the grey box of fig. 4. The first step involves identifying all humans in the scene using real-time RGB images from visual observations. Concurrently, a depth map, from the depth camera, estimates the distance of each detected human from the robot. If no human is detected within a predefined close range (e.g., 2 meters), the detection cycle repeats, and downstream HRC tasks are temporarily paused. This step ensures that the robot focuses on humans who are likely to interact closely, based on their proximity. Once a human is detected within the effective range, the RGB image is cropped to focus on the nearest individual. This cropped image is then processed by the pose estimation model to detect specific poses and key points. A sample outcome of this process is displayed in fig. 7. This hierarchical structure offers two main advantages: 1) By activating the pose estimation and planning modules only when a human is detected within a practical range, the system conserves computational and energy resources, operating more efficiently. 2) The hierarchical filtering effectively reduces the interference from non-target humans and background noise, thus improving the precision of pose detection. The following equation represents the operation of our proposed hierarchical pose detection model $\hat{f}_{\rm detection}$, which estimates the human pose $\hat{z}_{H}^{t}$ from visual observations: $\hat{z}_{H}^{t}=\hat{f}_{\rm detection}(O_{V}^{t})$.
IV-C Hierarchical Human Intention Prediction
Hierarchical human intention prediction is fundamental to effective long-term human-robot collaboration (HRC). This process involves two interconnected levels of intention analysis: human action prediction and human plan prediction. Action prediction focuses on low-level intentions that dictate short-term motion patterns, such as ”fetching a screw.” At a higher level, plan prediction addresses long-term intentions that are structured as sequences of such actions. As shown in the green box of fig. 4, the robot first predicts the human action from the current human pose. Then the high-level intention $\hat{x}_{H}^{t}$, or the human plan, considers both the current task progress and task graph.
IV-C1 Action Prediction
Figure 5: Action prediction pipeline

### v2

# Robustifying Long-term Human-Robot Collaboration through a Multimodal and Hierarchical Framework

@@ Abstract
Long-term Human-Robot Collaboration (HRC) is crucial for enabling flexible manufacturing systems and integrating companion robots into daily human environments over extended periods. This paper identifies several key challenges for such collaborations, such as accurate recognition of human plan, robustness to disturbances, operational efficiency, adaptability to diverse user behaviors, and sustained human satisfaction. To address these challenges, we model the long-term HRC task through a hierarchical task graph and presents a novel multimodal and hierarchical framework to enable robots to better assist humans to advance on the task graph. In particular, the proposed multimodal framework integrates visual observations with speech commands to facilitate intuitive and flexible human-robot interactions. Additionally, our hierarchical designs for both human pose detection and plan prediction allow better understanding of human behaviors and significantly enhance system accuracy, robustness and flexibility. Moreover, an online adaptation mechanism enables real-time adjustment to diverse user behaviors. We deploy the proposed framework to KINOVA GEN3 robot and conduct extensive user studies on real-world long-term HRC assembly scenarios. Experimental results show that our approaches reduce task completion time by 15.9%, achieves an average task success rate of 91.8% and an overall user satisfaction score of 84% in long-term HRC tasks, showcasing its applicability in enhancing real-world long-term HRC.
footnotetext: These authors contributed equally to this work.
@@ I INTRODUCTION
Human-robot collaboration (HRC) has become a critical area of research as robots emerge in human-centric environments. Significant progress has been made in short-term HRC [11, 4, 21], which addresses brief and localized interactions on single subtasks. On the other hand, long-term HRC studies extended and hierarchical collaboration involving multiple interdependent subtasks. For instance, in industrial settings, long-term HRC can significantly enhance productivity and adaptability in complicated manufacturing processes [32]. Similarly in home environments, robots acting as companions or elderly care agents can notably improve quality of life through long-term HRC [5]. Despite its importance in many applications, long-term HRC remains under-explored [6].
Achieving effective and robust long-term HRC presents several challenges: 1) robots must accurately interpret human actions and predict human plans in complex, extended tasks; 2) systems must remain robust against environmental noise and unexpected disturbances; 3) interactions should be efficient and optimized in completion time for varying contexts; 4) systems should adapt to diverse user behaviors and preferences; 5) systems should minimize human workload, ensure ease of use, and enhance satisfaction throughout the prolonged collaboration process.
Figure 1: Overview of the hierarchical task graph for the long-term toy car assembly HRC task. Each node in the long-term graph (left) corresponds to a short-term subtask (right), where the human and robot collaborate through sequential actions. The task consists of four stages and each stage is completed through a series of task nodes (e.g., must complete both Node 1 and Node 2 to complete The Bottom Stage ). The task starts at Node 1 and finishes at Node 7, with the directed arrows in the long-term task graph representing the temporal sequence of task execution (e.g., must complete Node 1 before performing Node 2). Both high-level task planning and short-term execution involve uncertainties. In the subtask graph, dashed lines indicate collaborative actions, while solid lines represent actions performed by a single agent. Arrows denote task progression. Arrows with a cross indicate robot motions that were initially planned but were not executed, as the robot adapted to human actions and switched to an alternative motion. For clarity, certain actions are grouped or omitted without affecting the task’s overall structure.
@@
Addressing these challenges is essential for deploying long-term HRC in real-world applications. Although existing research tackles some aspects of these challenges, there remains a critical need for integrated solutions that address several challenges simultaneously. For instance, studies on adaptable HRC [11, 4, 20] emphasize adaptable human motion prediction but often overlook robustness to disturbances or seamless integration required for effective HRC. By seamless integration, we refer to the robot’s ability to continuously interpret multimodal human inputs, adapt its behavior in real time, and transition smoothly between different subtasks without requiring frequent recalibration or explicit interventions [27]. Similarly, multimodal approaches to human plan understanding [19, 33] rely heavily on wearable devices, limiting their generalizability across different interaction contexts. Furthermore, many studies lack unified frameworks and often fail to demonstrate their applicability in real-world long-term HRC tasks [23].
In this work, we focus on enhancing long-term HRC characterized by extended, sequential interactions. We model long-term HRC tasks using a hierarchical task graph, assumed to be known to both human and robot participants. Fig. 1 illustrates an example of the task graph, consisting of a long-term task graph (left) and short-term subtask graphs (right). The long-term task graph structures the overall workflow, beginning at Node 1 with multiple pathways before converging at Node 7 (Assemble the Wheels) for task completion. The nodes are grouped into four stages. The short-term subtask graphs illustrate detailed human-robot interactions within each node, where sequential and collaborative actions are depicted.
The goal for the human and the robot is to progress from the initial node to the completion node of the task graph through a series of interactions. In this paper, we define a human plan as the route of nodes that the human expects to advance in order to complete the task. We divide the human plan into two levels: high-level plan (node sequences in the long-term graph) and low-level plan (node sequences in the short-term subtask graph). Given the long sequences of these interactions, long-term HRC is prone to compounding errors due to uncertainties in both high-level and low-level plans (e.g., the human can either go to node 3 or 5 from node 2 in the long-term task graph and may also have different low-level action plans in the sub-task graph). Therefore, precise and robust human plan predictions, and adaptability to individual users, are essential for effective long-term HRC.
To address these challenges, we propose a multimodal and hierarchical framework to enhance long-term HRC. The multimodality integrates visual and auditory signals, enabling accurate, reliable, and flexible predictions over dynamic and complex tasks. The hierarchical structure is reflected in two key components: hierarchical human pose detection and hierarchical plan prediction. In the plan prediction module, high-level human plans are first predicted based on the long-term task graph, and then refined through low-level plan predictions based on the subtask graph, improving flexibility to diverse user plans and enhancing system efficiency and task success rates. Hierarchical human pose detection further mitigates the effects of environmental noise and disturbances. Additionally, an adaptive action prediction module tailors the system to individual users. We apply the proposed multimodal hierarchical framework to a real-world human-robot co-assembly task. Experiment results demonstrate that our framework achieves precise, robust, flexible, and efficient collaboration, effectively addressing the critical challenges of long-term HRC.
@@
    We identify key challenges in long-term HRC tasks and propose a multimodal hierarchical framework that can effectively and robustly address these challenges.
@@
    We present a structured task graph for modeling long-term HRC tasks and deploy it to a real-world assembly scenario. Our proposed framework and task graph demonstrate successful HRC in the long-horizon assembly task.
@@
    We validate the framework through extensive physical experiments and user studies. The results show significant improvements in the robustness, time efficiency, and flexibility of HRC. Furthermore, user feedback highlights a strong preference for our framework, emphasizing its usability and practical applicability.
@@ II RELATED WORKS
Long-term Human-Robot Collaboration. Previous research on HRC has primarily focused on short-term tasks with long-horizon action sequences [28, 25], rather than fully addressing complex long-term tasks. Although these studies explore important aspects of task planning in extended action sequences [26, 14], there is limited work specifically targeting long-term HRC that achieves comprehensive and goal-oriented tasks. Pirk et al. [28] successfully recognized long-term human subassembly actions, but did not demonstrate HRC in real-world scenarios. In this work, we implement our framework on a long-term real-world human-robot co-assembly task.
@@
Human Pose Detection and Physical Action Prediction. High-fidelity and robust human behavior prediction is a key component of achieving safe HRC [2, 22]. Several approaches utilize human detection models, such as OpenPose [9], to identify keypoints on human body, which are then used to predict human behaviors including intentions and/or trajectories [4, 20]. Reducing disturbances in human pose detection has been extensively studied using signal filtering techniques [34, 1]. Traditional filtering methods primarily mitigate temporal disturbances [30]. In contrast, we focus on spatial aspects and propose a novel hierarchical framework to reduce disturbances in HRC. Additionally, we implemented an online model adaptation algorithm to adapt the prediction models to accommodate different users’ behaviors [3].
@@ III PROBLEM FORMULATION
Long-Term HRC Task. This paper focuses on assistive collaboration in the human-robot supervisor-subordinate paradigm [18]. This paradigm leverages human strengths while enabling robots to reduce the physical and cognitive demands on humans by handling repetitive, heavy, or precision-intensive tasks. This approach is particularly well-suited for human-robot collaboration in manufacturing settings, where efficiency, workload distribution, and adaptability are critical.
@@
We consider a scenario where a human $A_{H}$ collaborates with a robot $A_{R}$, to achieve a set of predefined goals within a long-term task $\mathcal{G}$. The task $\mathcal{G}$ is represented as a hierarchical directional And-Or graph [15], with the ultimate goal $g$ defined as reaching the end node of the graph (see Fig. 1 left). The nodes $\mathcal{N}$ of $\mathcal{G}$ correspond to subtasks, and the directed edges represent precedence relationships. We group several nodes into one stage, and each stage may have multiple successors, reflecting the possibility of alternative plans (e.g., ”The Bottom” stage in Fig. 1 can lead to either ”The Middle” or ”The Top” stage).
Long-term HRC tasks pose significant challenges due to uncertainties and disturbances, particularly in human behavior. In this work, we focus on human-related uncertainties while assuming object positions are well-observed, as human behavior and actions present greater unpredictability. Key challenges include variability (including diverse preferences and styles) in human behavior, visual ambiguities caused by occlusions or background clutter, and uncertainty in human plans at both high and low levels of the task graph. Additionally, environmental disturbances and compounding errors over prolonged interactions exacerbate these challenges. To address these challenges, we aim to robustify long-term HRC by focusing on human behavior-related uncertainties and disturbances.
@@
Sub-task Configuration. Nodes $\mathcal{N}$ in the long-term task $\mathcal{G}$ corresponds to shorter-term HRC tasks, referred to here as subtasks. As shown in the right part of Fig. 1, each subtask $N_{i}\in\mathcal{N}$ is organized as a multi-agent Temporal Plan Graph (TPG), which represents both the precedence and collaborative dynamics between human and robot actions [13]. A Temporal Plan Graph (TPG) [17] captures precedence relationships within a HRC and preserves them during execution. Each vertex in $N_{i}$ represents an action performed by a human or a robot.
Definition of Hierarchical Human Plan. The human plan refers to the route of nodes that the human is expecting to advance on to complete the task. We divide human plan $x_{H}$ into two levels, where the low-level plan is a sequence of nodes in a short-term sub-task graph and the high-level plan is a sequence of nodes in a long-term task graph. Fig. 1 illustrates that the user may have different high-level plans as well as different low-level plans within the same high-level plan. To complete the long-term HRC task, the robot must predict both human plans precisely and perform corresponding responses.
@@
The Objective Function. Let $E$ denote the environment state in the HRC task, including human actions, object positions, and surrounding context. $E$ is derived from sensor observations, which may include noise or ambiguities in human behavior. Let $P$ denote the task progress, which tracks the completion of subtasks (or the current node within the task graph $\mathcal{G}$) by incorporating both the sequence of human-robot actions executed and the overall task graph. As the collaboration progresses, $P$ is dynamically updated. At timestep $t$, the human’s plan $x_{H}^{t}$ is influenced by the current state of the environment $E^{t}$ and task progress $P^{t}$. The robot predicts human plans $\hat{x}_{H}^{t}$ using its learnable model $\hat{f}_{H}$:
@@
The predicted plans guide the robot in generating a motion sequence $q=[q^{1},\cdots,q^{n}]$. The objective for the robot is to find an efficient and robust motion sequence $q$ under safety and dynamics constraints. Here, $n$ represents the planning horizon, which is determined based on the specific task, the robot’s capabilities, and the reliability of the predictions. For example, $n$ may correspond to planning the robot’s motion for the next 2 seconds based on the current predicted human plans. Let $J_{g}$ represent the task-specific cost of achieving the goal $g$, such as the tracking cost. To reflect time efficiency, we include the task completion time $T_{g}$, weighted by $\lambda$, in the objective function. Thus, the robot’s objective in HRC is formulated as:
@@
where $M_{R}(\cdot)$ represents the robot motion planning function. This objective is influenced by the environment state $E^{t}$, task progress $P^{t}$, and predicted human plan $\hat{x}_{H}^{t}$. While $E^{t}$ and $P^{t}$ are relatively deterministic, the uncertainty in human plans and variability in human behavior significantly impact the generation of optimal robot motions. Thus, robust and accurate prediction of human plans, along with mitigating the effects of noise on human behavior, is crucial for achieving optimal motion generation. This, in turn, is essential for efficient and robust HRC.
@@ IV PROPOSED FRAMEWORK
To address the challenges of long-term HRC tasks as discussed in Sec. I, we propose a hierarchical and multimodal HRC framework, as illustrated in Fig. 2. The framework includes two perception modules, four task planning modules, and a robot motion controller.
@@
  • Planning. Human Pose Detection Module identifies human poses and extracts key point trajectories. Its hierarchy enhances detection robustness by mitigating disturbances in complex, multi-human environments. Human Physical Action Prediction Module predicted future human pose trajectories and interpret the semantic action label of the current human pose. Human Speech Action Recognition Module recognizes human speech commands and outputs human speech action labels. Human Plan Prediction Module integrates physical and speech actions and task progress to predict human plans. Its hierarchical design improves prediction accuracy and enhances system robustness in responding to human actions [11].
  • Control. Controller Module synthesizes predicted human plans and advances on the task graph by generating executable robot motions.
@@ IV-A Multimodal Perception
In human-human interactions, visual observations alone often fail to fully capture human plans, necessitating the integration of verbal communication for comprehensive context. Extending this principle to long-term HRC tasks, we propose a multimodal framework (depicted in Fig. 2) that combines visual and audio observations for more comprehensive perception. As the same human physical action may indicate different plans depending on the context, we incorporate speech actions to enable humans to clarify and refine robot-predicted human plans derived from visual interpretations. The multimodal approach offers two key advantages: 1) By integrating visual and audio signals, the robot achieves a comprehensive perception of human plans, allowing for more complex and long-term collaborations. 2) The inclusion of speech enables more natural and flexible interactions, making the system more user-friendly and accessible.
Let $O_{V}$ denote the robot’s visual observation from the environment $E$, captured by the RGBD camera; and $O_{A}$ denote the robot’s audio observation, captured by the microphone. It can be easily deducted that multual information $I(g;O_{V},O_{A})\geq\max\{I(g;O_{V}),I(g;O_{A})$}. The proof of this inequality is provided in the appendix. This inequality demonstrates that the multimodal framework yields greater mutual information than single-modal approaches, significantly enhancing the robot’s ability to manage complex and long-term collaborative tasks.
IV-B Hierarchical Human Pose Detection
Human poses are fundamental to understanding human physical actions, which in turn are critical for predicting human plans. However, visual observations can often be noisy, leading to errors in pose detection and subsequent plan prediction. This noise can arise from various sources, such as the presence of multiple humans in the scene or cluttered backgrounds. To mitigate these disturbances, we propose a hierarchical pose detection module.
Our module operates in two steps, as illustrated in the yellow box in Fig. 2. First, we detect all humans in the scene and estimate their distances from the robot through an RGB-D camera. If no human is detected within a predefined range (e.g., 2 meters), the detection cycle repeats, and downstream HRC tasks are temporarily paused. This ensures the robot focuses on humans likely to interact closely based on proximity. Once a human is detected within the effective range, we crop on the closest human. This cropped image is then processed by the pose estimation model to extract its key points. This design significantly reduces keypoint deviation, thus help to make pose detection more robust, as shown in Fig. 6.
This hierarchical structure offers two key advantages: 1) By activating pose estimation and planning modules only when a human is detected within range, the system conserves computational and energy resources. 2) The hierarchical filtering reduces interference from non-target humans and background noise, improving pose detection accuracy. The following equation represents the operation of our proposed hierarchical pose detection model $\hat{f}_{\rm detection}$, which estimates the human pose $\hat{z}_{H}^{t}$ from visual observations: $\hat{z}_{H}^{t}=\hat{f}_{\rm detection}(O_{V}^{t})$.
IV-C Human Physical Action Prediction